OpenStudy (anonymous):

Simplify the expression using the reciprocal relations. 1. cos theta sec theta tan theta a. sin theta b. 1 c. 0 d. tan theta e. cos theta

OpenStudy (kropot72):

\[\cos \theta \times \sec \theta \times \tan \theta\] \[\sec \theta=\frac{1}{\cos \theta}\] Can you find the answer by substituting the reciprocal relation for sec theta?

OpenStudy (anonymous):

No, im not sure how to do that.

OpenStudy (kropot72):

\[\frac{\cos \theta \times \tan \theta}{\cos \theta}=?\]

OpenStudy (anonymous):

tan theta?

OpenStudy (kropot72):

tan theta is the right choice :)

OpenStudy (anonymous):

Yayy! (:

OpenStudy (anonymous):

can you help me with 4 more?

OpenStudy (kropot72):

Sorry I can't promise to be available to give more help right now>

OpenStudy (anonymous):

Ok, well thank you.